OpenStudy (anonymous):

need help finding the inverse of this equation: sqrt ( x^2 +5x)

OpenStudy (anonymous):

\[\sqrt{x^2+5x}\] need to find inverse of that

OpenStudy (anonymous):

so first I do x= \[\sqrt{y^2+5y}\]

OpenStudy (anonymous):

then \[x^2 = y^2 + 5y\]

OpenStudy (anonymous):

but then I get confused about what to do next

OpenStudy (anonymous):

yeah..i really dont know how to do this lol.

OpenStudy (anonymous):

yeh thats ur inverse u got it man

OpenStudy (anonymous):

But that's not the answer my textbook has at all. My textbook somehow gets it down to a y = equation

OpenStudy (anonymous):

it's solution says \[-\frac{ 5 }{ 2 } + \sqrt{x^2 + (\frac{ 5 }{ 2 })^2}\]

OpenStudy (anonymous):

both are same actually